OpenStudy (anonymous):

Inner product space question: I have trouble understanding this equality. ||^2=conj() Where |z| is absolute value of z x,y are vectors from the inner product space is the inner product of vectors x and y conj(z) is the complex conjugate of z

OpenStudy (phi):

What exactly is the question? do you agree that the magnitude of a complex number z= a+b i is |z|= sqrt(a^2+b^2) or is that your question?

OpenStudy (anonymous):

Ah, I think I'm getting there, thanks for replying! (Although if you'd feel like clarifying more, I wouldn't protest) I haven't done any complex analysis yet, so those parts of linear algebra where understanding of complex numbers is required often confuse me. Thanks.

OpenStudy (anonymous):

Okay, a few minutes of math paperwork and I've got it. I really ought to study a bit of complex analysis!

OpenStudy (phi):

one way to see that the magnitude of z squared is a^2+b^2: